End of Section Problem 9.4 Use the data given to test the following hypotheses. Assume the data are normally distributed in the population. Ho : ?-7.48 HH<7.48 Tab (Round your answer to 2 decimal places.) The value of the test statistic is and we
Answer #1

Solution :

Given that,

\bar x = 6.91

\mu = 7.48

\sigma = 1.21 and n = 24

z = (\bar x - \mu) / \sigma / \sqrtn = (6.91 - 7.48) / 1.21 / \sqrt24 = -2.31

This is the left tailed test .

P(z < -2.31) = P(z < -2.31) = 0.0104

Here , P-value > \alpha

We do not reject the null hypothesis .

P-value = p = 0.0301